# Spectroscopy and complex-time correlations using minimally entangled typical thermal states

### v1

@@ I Introduction
As we will show, complex-time correlation functions can be evaluated using well-known MPS time evolution algorithms such as the time-dependent variation principle (TDVP) [54, 55]. Our aim is to combine the advantages of computing imaginary-time correlations and real-time correlations. The imaginary contribution of time evolution dissipates the large entanglement entropy while the real-time part of the correlation function benefits the efficiency of obtaining real frequency spectral function.
@@
The outline of our paper is as follows. We begin with Sec. II introducing the Shastry-Sutherland model we use to benchmark out methods. In Sec. III.1, we review the conventional thermal METTS algorithm of simulating static observables at nonzero temperatures. The dynamical METTS algorithm of obtaining spectral function via real-time correlations is introduced in Sec. III.2. As both these algorithms are based on efficient matrix-product state time-evolution algorithms, we discuss the time-evolution methods employed in Sec. III.3. Details on the Fourier analysis of time-dependent correlation functions are discussed in Sec. III.4. To validate our approach, we perform a comparison of the dynamical METTS algorithm to spectral functions obtained via the Lehmann representation using full exact diagonalization and DMRG in Sec. III.5.
@@
where $\bm{r}_{i}$ denotes the position of site $i$ and $N$ denotes the number of lattice sites. The position of the lattice sites $\bm{r}_{i}$ as well as the unit cell and basis sites are considered in the square lattice geometry shown in Fig. 1(b). To reduce the boundary effects of the cylinder we use a modified definition,
@@ III.1 Review of the thermal METTS algorithm
where $p_{i}=\langle\sigma_{i}|e^{-\beta H}|\sigma_{i}\rangle\geq 0$ is is a (non-negative) weight. Notice that the partition function can also be written as $\mathcal{Z}=\sum_{i}p_{i}$. The METTS states are normalized, i.e., $|\langle\psi_{i}|\psi_{i}\rangle|^{2}=1$. To sample the METTS states $|\psi_{i}\rangle$ with probability $p_{i}/\mathcal{Z}$ a Markov chain is constructed where the transition probability is given by,
@@
and, therefore, the stationary distribution of this Markov chain is given by the probabilities $p_{i}$. As a Markov chain Monte Carlo simulation, thermalization and autocorrelation effects have to be taken into account to estimate averages and standard errors. The METTS algorithm is particularly well-suited to tensor network simulations describing states of low entanglement. Due to the fact that we sample with non-negative probabilities $p_{i}$ the Monte Carlo sampling itself is not accompanied by a sign problem, as often encountered in QMC simulations. The computational complexity is rather transferred to the entanglement of the METTS states |$\ket{\psi_{i}}$.
@@ III.2 Dynamical METTS algorithm in real-time
Figure 3: Exemplary METTS samples (blue lines) of (a) the correlation functions $\mathcal{C}_{AB}^{i}(t)$ and (b) the spectral functions $\mathcal{C}_{AB}^{i}(\omega)$ of the Shastry-Sutherland model on a $16\times 4$ cylinder for $J/J_{D}=0.63$ and $T/J_{D}=0.1$. We investigate the spin correlations and structure factors with $A^{\dagger}=B=S^{z}(\bm{q})$ as in Eqs. 3 and 4 with $\bm{q}=(\pi,0)$. The estimated averages $\mathcal{C}_{AB}(t)$ and $\mathcal{C}_{AB}(\omega)$ with a statistical error estimate from $R=100$ samples are shown as the red line with the shaded light red region.
@@
as defined in Eqs. 3 and 4. A dynamical spectral function is then given as the Fourier transform,
@@
     Compute the METTS state $|\psi_{i}\rangle=\frac{1}{\sqrt{p_{i}}}e^{-\beta H/2}|\sigma_{i}\rangle$
@@
Algorithm 1 Dynamical METTS for correlators $\mathcal{C}(t)$
The next step is to perform a real-time evolution to compute $|v_{i}(t)\rangle$ and $|w_{i}(t)\rangle$ as in Eqs. 17 and 18. Notice that the METTS states $|\psi_{i}\rangle$ are not eigenstates of the Hamiltonian $H$, such that the time-evolved states $|v_{i}(t)\rangle$ and $|w_{i}(t)\rangle$ have to be explicitly computed. Naturally, this time evolution can only be carried out until a maximum time $\Omega$. Measurements of $\mathcal{C}_{AB}^{i}(t)$ are only performed at discrete time steps which are a multiple of an elementary step size $\delta_{t}$.
@@
The algorithm laid out above is summarized in Alg. 1, and illustrated graphically in Fig. 2. Using the Shastry-Sutherland model as a concrete example, we show the corresponding correlation function $\mathcal{C}_{AB}^{i}(t)$ and $\mathcal{C}_{AB}^{i}(\omega)$ for several sampled METTS configurations as well as their average value in Fig. 3.
The real-time approach of measuring spectral functions at finite temperature is rigorous as in the limit of long time domain and sufficient sampling of METTS states |$\ket{\psi_{i}}$. However, a limitation of this algorithm is naturally the unbounded growth of entanglement entropy in real-time evolution. In particular, at higher temperatures entanglement growth can hinder accurate calculations at given computational resources. As shown in Fig. 6 below, the bipartite entanglement entropy $S_{\text{vN}}$ of time evolution based on METTS states of the Shastry-Sutherland model at $T/J_{D}=0.2$ grows much more quickly on average than the one at $T/J_{D}=0.1$. Meanwhile, pronounced fluctuations of the entropy after time evolution exist among various sampled METTS configurations.
@@ III.5 Validation by Exact Diagonalization and DMRG
To address these questions, we show results from ED as well as dynamical METTS simulations in Fig. 4. We focus on the parameter $J/J_{D}=0.63$ and $\bm{q}=(\pi/2,\pi/2)$ and consider two temperatures $T/J_{D}=0.10$ in (a) and $T/J_{D}=0.20$ (b). We sampled $1000$ METTS and performed a time-evolution up to a maximal time of $\Omega/J_{D}=50$ for every $10$-th METTS. By only considering a subset of all sampled METTS, we verified that the measurement time series has a negligible autocorrelation time. We also employed measurements in the $S^{x}$-basis, to further assure vanishing autocorrelation times, see e.g. [13, 14]. We employed 2TDVP as the time evolution algorithm with a cutoff $\varepsilon=10^{-6}$ without a constraint on the maximal bond dimension $\chi_{\max}$. While a $4\times 4$ simulation cluster is arguably small, we would like to point out that the employed fully periodic boundary conditions are not without challenge for MPS methods, since long-range interactions are present. Hence, this benchmark also serves as a good first test of the quality of time-evolution algorithms employed to compute the METTS states as well as the real-time evolution. To compute the Fourier transform $S(\bm{q},\omega)$ to frequency space, we compared different windowing functions and found the Hann window Eq. 30 to give the most accurate results, cf. Sec. III.4. For ED, we compare three different values of the broadening $\eta=0.02,0.05,0.10$.
@@
We report a close agreement between the ED and METTS simulations. While both the Gaussian broadening $\eta$ for the ED data as well as the finite time horizon $\Omega$ have a broadening effect on the exact spectrum Eq. 33, we observe that even small intricate structures of the spectrum (likely stemming from the small system size) are accurately captured by the METTS algorithm. In particular, the positions of the numerous side peaks are accurately reproduced, as well as the overall spectral amplitude. Choosing $\Omega/J_{D}=50$ and applying the windowing procedure as explained in Sec. III.4 yields a broadening of the spectral function comparable to a Gaussian broadening $\eta\approx 0.05$ in ED. Hence, we conclude that having a finite time horizon $\Omega$ as well as only a discrete set of measurements with time steps $\delta_{t}$ only leads to a small broadening of the spectrum. Moreover, we report that averaging over only $R=100$ METTS yields a satisfyingly small statistical error indicated by the red-shaded region.
@@
We now compare the results from simulations at nonzero temperature with ground state spectral functions from DMRG. This demonstrates the accuracy of the dynamical METTS algorithm in the limit of low temperatures. Results for spectral functions at different temperatures, including $T=0$ from DMRG, are shown in Fig. 5. We focus on the $16\times 4$ cylinder at $J/J_{D}=0.63$ and two wave vectors $\bm{q}=(\pi,0)$ and $\bm{q}=(\pi/2,\pi/2)$. Our maximal time horizon for the real-time evolution has been chosen as $\Omega/J_{D}=50$. Moreover, we collected $R=100$ independent METTS samples for each temperature and observed that this allows sharp error estimates with a large signal-to-noise ratio, as apparent by the small error bars in Fig. 5. The time evolution has been performed using 2TDVP with maximal bond dimensions $\chi_{\max}=1000,1500,2000$. Comparing different bond dimensions for different seeds leads us to conclude that for this particular system, the sample real-time correlation functions $C^{i}(t)$ are already well-converged at $\chi_{\max}=1000$. The spectral function from DMRG has similarly been obtained by a single time-evolution up to time $\Omega/J_{D}=50$. We observe a structured spectrum, where the dominant peak corresponds to single triplon excitations, but also secondary and tertiary peaks corresponding to multi-triplon excitations of the Shastry-Sutherland model. The good convergence towards the DMRG results for $T\rightarrow 0$, where even small features such as secondary and tertiary peaks are well-resolved, demonstrates the accuracy of the approach at low temperatures.
@@ IV.1 Two variants of complex-time correlation functions
Real-time evolution of MPS is computationally expensive due to unbounded entanglement growth for generic wave functions. Imaginary-time correlation functions, on the other hand, can be efficiently computed since for imaginary-time $\tau\rightarrow\infty$ every wavefunction $|\phi\rangle$ will converge to the ground state $|\phi\rangle=|\psi_{0}\rangle$ whenever both wave functions have non-vanishing overlap $|\langle\phi|\psi_{0}\rangle|\neq 0$, i.e.
@@
There are two conceivable alternatives when defining complex time correlation functions,
@@
The hermitian correlator $\mathcal{H}_{AB}(z)$ as in Eq. 43, on the other hand, has only recently been investigated in the context of MPS and impurity models [53] and has several intriguing properties which are discussed Sec. IV.4. We choose the name hermitian, since whenever $A$ is a hermitian operator, i.e., $A^{\dagger}=A$, the time-evolved operator $\mathcal{U}(z)^{\dagger}A\mathcal{U}(z)$ is also hermitian. There are computational advantages in evaluating this time-dependent correlation function. For convenience, we again introduce the shorthand notations,
@@
Figure 7: Sketch of the complex contour on which the analytical correlation function $\mathcal{A}(z)$ is evaluated. The maximal imaginary time is chosen to be $\Omega_{\tau}=\beta/2$ in order to allow for an efficient and stable numerical algorithm.
@@ IV.2 Analytic complex-time correlation functions
$\mathcal{A}(z)$ $=\braket{e^{iHz}Ae^{-iHz}B}_{\beta}$ (47)
@@
$|v_{i}(t)\rangle=e^{-iHz}B|\psi_{i}\rangle=e^{-iHt}e^{-\tau H}B|\psi_{i}\rangle$ (55)
$|w_{i}(t)\rangle=e^{-iH\overline{z}}|\psi_{i}\rangle=e^{-iHt}e^{+\tau H}|\psi_{i}\rangle,$ (56)
with $z=t-i\tau$ similar to the real-time correlation algorithm Alg. 1. However, this approach is plagued by severe numerical instability when computing the state $|w_{i}(t)\rangle$. Computing a positive exponential $e^{+\tau H}|\psi_{i}\rangle$ is numerically highly unstable and prohibits most practical calculations.
When restricting our time horizon to $\Omega_{\tau}=\beta/2$ we can circumvent ever having to compute positive real exponential in the context of the METTS algorithm. The key idea is that the positive imaginary time evolution can be absorbed by the negative imaginary time evolution performed when calculating an individual METTS |$\ket{\psi_{i}}$. For this, we rewrite Eq. 47 as,
Algorithm 2 Dynamical METTS for analytic correlator $\mathcal{A}(z)$
@@
$\mathcal{A}$ $(z)=\braket{e^{izH}Ae^{-izH}B}_{\beta}=$ (57)
$=\braket{e^{-i\Omega_{t}H}e^{izH}Ae^{-izH}Be^{i\Omega_{t}H}}_{\beta}$ (58)
$=\frac{1}{\mathcal{Z}}\sum_{i}p_{i}\langle\psi_{i}|e^{-i\Omega_{t}H}e^{izH}Ae^{-izH}Be^{i\Omega_{t}H}|\psi_{i}\rangle$ (59)
$=\frac{1}{\mathcal{Z}}\sum_{i}p_{i}\;\mathcal{A}^{i}(z),$ (60)
where we defined the sample analytic correlation function as,
$\mathcal{A}^{i}(z)$ $=\langle\psi_{i}|e^{-i\Omega_{t}H}e^{izH}Ae^{-izH}Be^{i\Omega_{t}H}|\psi_{i}\rangle$ (62)
Using a discretized $z=k\delta_{z}$ we can rewrite Eq. 63 as,
$\mathcal{A}^{i}(k\delta_{z})=\bra{\phi^{S-k}_{i}}A\ket{\chi_{i}^{k}}.$ (68)
@@
To evaluate these expectation values for all discretized values of $z=k\delta_{z}$ ($k=0,\ldots,S$), we compute the states |$\ket{\phi_{i}^{k}}$ for $k=0,\ldots,S$ successively by complex time evolution and store all intermediate states to disk for later use. Then we compute the state |$\ket{\chi^{0}_{i}}=B\ket{\psi_{i}^{S}}$ from which we successively compute |$\ket{\chi^{k}_{i}}$ for $k=1,\ldots,S$ and thereby evaluate the desired correlation functions $\mathcal{A}^{i}(z)=\bra{\phi^{S-k}_{i}}A\ket{\chi_{i}^{k}}$. $\mathcal{A}(z)$ is then obtained by averaging over $\mathcal{A}^{i}(z)$. We summarize the described algorithm in Alg. 2.
@@
Alg. 2 is in several aspects similar to Alg. 1. Both algorithms require computing the METTS |$\ket{\psi_{i}}$ and then time-evolves two further states to compute time-dependent correlators. Moreover, the basic thermal METTS Markov chain is constructed in both cases. A difference in Alg. 2 is the necessity of storing the auxiliary states |$\ket{\phi_{i}^{k}}$. On present-day computers, this typically neither imposes severe time nor memory constraints. However, in the limit of small angle $\theta$ the memory requirement can be considerable due to the large number of discrete time steps and requires corresponding storage capacities.
The key quantity determining the efficiency of the algorithm is the bipartite von-Neumann entanglement entropy $S_{\text{vN}}$ as in Eq. 35 of the MPS |$\ket{\phi_{i}^{k}}$ and |$\ket{\chi_{i}^{k}}$, shown in Figs. 8 and 9, respectively. The entanglement is considered for a bipartition at the center of the system. We evaluated the analytic correlator for the dynamical spin structure factor $S^{z}(\bm{q},\omega)$ as in Eq. 2 for the $16\times 4$ Shastry-Sutherland model at $J/J_{D}=0.63$ at temperatures $T/J_{D}=0.1$ and $T/J_{D}=0.2$. The auxiliary state |$\ket{\phi_{i}^{0}}=\ket{\sigma_{i}}$ is a classical product state such that the entropy in Fig. 8 starts from zero. Slower entanglement growth is observed for larger values of $\theta$ as expected. For $\theta=\pi/2$ we are performing fully imaginary time evolution, which yields the slowest entanglement growth. Meanwhile, fluctuations of $S_{\text{vN}}$ among the sampled states during METTS dynamics are rather small, as indicated by the shaded regions around the mean. The absolute values of $S_{\text{vN}}$ are comparable for both |$\ket{\phi_{i}^{k}}$ and |$\ket{\chi_{i}^{k}}$ and grow with increasing temperature. We observe a significant reduction in entanglement entropy when comparing the values to the entanglement of the states $|v_{i}(t)\rangle$ Eq. 17 as shown in Fig. 6. While we observed maximal values of $S_{\text{vN}}\approx 5.3$ for $T/J_{D}=0.2$ for $|v_{i}(t)\rangle$ in Fig. 6, we only find maximal values of $S_{\text{vN}}\approx 2.3$ for $\theta=\pi/8$ in Figs. 8 and 9. Naturally, the chosen time horizon $\Omega$ is significantly shorter when evaluating the analytical correlator. However, this demonstrates that $\mathcal{A}(z)$ can be evaluated at significantly reduced cost at non-zero values of $\theta$. We want to point out that setting a value of $\theta$ determines the value of $\Omega$ due to the constraint $\Omega_{\tau}=\beta/2$ we impose for the analytical correlator. Hence, the maximal complex times $|z|$ differ for different values of $\theta$.

### v2

@@ I Introduction
As we will show, complex-time correlation functions can be evaluated using well-known MPS time evolution algorithms such as the time-dependent variational principle (TDVP) [54, 55]. Our aim is to combine the advantages of computing imaginary-time correlations and real-time correlations. The imaginary contribution of time evolution dissipates the large entanglement entropy while the real-time part of the correlation function benefits the efficiency of obtaining real frequency spectral function.
@@
[…]
The outline of our paper is as follows. We begin with Sec. II introducing the Shastry-Sutherland model we use to benchmark our methods. In Sec. III.1, we review the conventional thermal METTS algorithm of simulating static observables at nonzero temperatures. The dynamical METTS algorithm of obtaining spectral function via real-time correlations is introduced in Sec. III.2. As both these algorithms are based on efficient matrix-product state time-evolution algorithms, we discuss the time-evolution methods employed in Sec. III.3. Details on the Fourier analysis of time-dependent correlation functions are discussed in Sec. III.4. To validate our approach, we perform a comparison of the dynamical METTS algorithm to spectral functions obtained via the Lehmann representation using full exact diagonalization and DMRG in Sec. III.5.
@@
where $\bm{r}_{i}$ denotes the position of site $i$ and $N$ denotes the number of lattice sites. The position of the lattice sites $\bm{r}_{i}$ as well as the unit cell and basis sites are considered in the square lattice geometry shown in Fig. 1(b). he Shastry-Sutherland model features a flat band of triplon excitations, which stretches throughout the entire Brillouin zone. At low temperatures we expect similar physical phenomena to occur regardless of the exact momentum. In this manuscript we focus on two momenta, $\bm{q}=(\pi,0)$ and $\bm{q}=(\pi/2,\pi/2)$ to illustrate the physics, but expect similar behavior at different values of $\bm{q}$.
To reduce the boundary effects of the cylinder we use a modified definition,
@@ III.1 Review of the thermal METTS algorithm
where $p_{i}=\langle\sigma_{i}|e^{-\beta H}|\sigma_{i}\rangle\geq 0$ is a (non-negative) weight. Notice that the partition function can also be written as $\mathcal{Z}=\sum_{i}p_{i}$. The METTS states are normalized, i.e., $|\langle\psi_{i}|\psi_{i}\rangle|^{2}=1$. To sample the METTS states $|\psi_{i}\rangle$ with probability $p_{i}/\mathcal{Z}$ a Markov chain is constructed where the transition probability is given by,
@@
and, therefore, the stationary distribution of this Markov chain is given by the probabilities $p_{i}$. As a Markov chain Monte Carlo simulation, thermalization and autocorrelation effects have to be taken into account to estimate averages and standard errors. In the following, we denote by $R$ the number of random METTS states sampled. In our calculations, we choose $R=100$ due to our computational budget, from which we obtain error estimates by analyzing thermalization and autocorrelation times as in standard error analysis of Markov chain Monte Carlo methods. The METTS algorithm is particularly well-suited to tensor network simulations describing states of low entanglement. Due to the fact that we sample with non-negative probabilities $p_{i}$ the Monte Carlo sampling itself is not accompanied by a sign problem, as often encountered in QMC simulations. The computational complexity is rather transferred to the entanglement of the METTS states |$\ket{\psi_{i}}$.
@@ III.2 Dynamical METTS algorithm in real-time
as defined in Eqs. 4 and 5. A dynamical spectral function is then given as the Fourier transform,
@@
     1 Compute the METTS state $|\psi_{i}\rangle=\frac{1}{\sqrt{p_{i}}}e^{-\beta H/2}|\sigma_{i}\rangle$
@@
Dynamical METTS for correlators $\mathcal{C}(t)$
where,
[…]
The next step is to perform a real-time evolution to compute $|v_{i}(t)\rangle$ and $|w_{i}(t)\rangle$ as in Eqs. 18 and 19. Notice that the METTS states $|\psi_{i}\rangle$ are not eigenstates of the Hamiltonian $H$, such that the time-evolved states $|v_{i}(t)\rangle$ and $|w_{i}(t)\rangle$ have to be explicitly computed. Naturally, this time evolution can only be carried out until a maximum time $\Omega$. Measurements of $\mathcal{C}_{AB}^{i}(t)$ are only performed at discrete time steps which are a multiple of an elementary step size $\delta_{t}$.
Figure 3: Exemplary METTS samples (blue lines) of (a) the correlation functions $\mathcal{C}_{AB}^{i}(t)$ and (b) the spectral functions $\mathcal{C}_{AB}^{i}(\omega)$ of the Shastry-Sutherland model on a $16\times 4$ cylinder for $J/J_{D}=0.63$ and $T/J_{D}=0.1$. We investigate the spin correlations and structure factors with $A^{\dagger}=B=S^{z}(\bm{q})$ as in Eqs. 4 and 5 with $\bm{q}=(\pi,0)$. The estimated averages $\mathcal{C}_{AB}(t)$ and $\mathcal{C}_{AB}(\omega)$ with a statistical error estimate from $R=100$ samples are shown as the red line with the shaded light red region.
@@
The algorithm laid out above is summarized in Fig. 2, and illustrated graphically in Fig. 2. Using the Shastry-Sutherland model as a concrete example, we show the corresponding correlation function $\mathcal{C}_{AB}^{i}(t)$ and $\mathcal{C}_{AB}^{i}(\omega)$ for several sampled METTS states as well as their average value in Fig. 3. At the given temperature $T/J_{D}=0.1$ we observe a “bunching” of correlation functions $\mathcal{C}_{AB}^{i}(t)$ around the $T=0$ correlation function whereas several other random realizations disperse, illustrating the effect of thermal fluctuations.
The real-time approach of measuring spectral functions at finite temperature is rigorous as in the limit of long time domain and sufficient sampling of METTS states |$\ket{\psi_{i}}$. However, a limitation of this algorithm is naturally the unbounded growth of entanglement entropy in real-time evolution. In particular, at higher temperatures entanglement growth can hinder accurate calculations at given computational resources. As shown in Fig. 6 below, the bipartite entanglement entropy $S_{\text{vN}}$ of time evolution based on METTS states of the Shastry-Sutherland model at $T/J_{D}=0.2$ grows much more quickly on average than the one at $T/J_{D}=0.1$. Meanwhile, pronounced fluctuations of the entropy after time evolution exist among various sampled METTS states.
@@ III.5 Validation by Exact Diagonalization and DMRG
To address these questions, we show results from ED as well as dynamical METTS simulations in Fig. 4. We focus on the parameter $J/J_{D}=0.63$ and $\bm{q}=(\pi/2,\pi/2)$ and consider two temperatures $T/J_{D}=0.10$ in (a) and $T/J_{D}=0.20$ (b). We sampled $1000$ METTS and performed a time-evolution up to a maximal time of $\Omega/J_{D}=50$ for every $10$-th METTS. By only considering a subset of all sampled METTS, we verified that the measurement time series has a negligible autocorrelation time. We also employed measurements in the $S^{x}$-basis, to further assure vanishing autocorrelation times, see e.g. [13, 14]. We employed 2TDVP as the time evolution algorithm with a cutoff $\varepsilon=10^{-6}$ without a constraint on the maximal bond dimension $\chi_{\max}$. While a $4\times 4$ simulation cluster is arguably small, we would like to point out that the employed fully periodic boundary conditions are not without challenge for MPS methods, since long-range interactions are present. Hence, this benchmark also serves as a good first test of the quality of time-evolution algorithms employed to compute the METTS states as well as the real-time evolution. We apply the MPO representing $S^{z}(\bm{q})$ directly for a specific $q$ in our simulations instead of measuring correlations between pairs of sites. This is done in order to minimize the amount of measurements needed, and to allow for statistical self-averaging across the lattice during sampling. To compute the Fourier transform $S(\bm{q},\omega)$ to frequency space, we compared different windowing functions and found the Hann window Eq. 31 to give the most accurate results, cf. Sec. III.4 and appendix C. For ED, we compare three different values of the broadening $\eta=0.02,0.05,0.10$.
@@
We report a close agreement between the ED and METTS simulations. While both the Gaussian broadening $\eta$ for the ED data as well as the finite time horizon $\Omega$ have a broadening effect on the exact spectrum Eq. 34, we observe that even small intricate structures of the spectrum (likely stemming from the small system size) are accurately captured by the METTS algorithm. In particular, the positions of the numerous side peaks are accurately reproduced, as well as the overall spectral amplitude. Choosing $\Omega/J_{D}=50$ and applying the windowing procedure as explained in Sec. III.4 yields a broadening of the spectral function comparable to a Gaussian broadening $\eta\approx 0.05$ in ED. Hence, we conclude that having a finite time horizon $\Omega$ as well as only a discrete set of measurements with time steps $\delta_{t}$ only leads to a small broadening of the spectrum. Discrepancies around the $\omega\rightarrow 0$ limit stem from the finite simulation time-horizon $\Omega$. Moreover, we report that averaging over only $R=100$ METTS yields a satisfyingly small statistical error indicated by the shaded regions.
@@
We now compare the results from simulations at nonzero temperature with ground state spectral functions from DMRG. This demonstrates the accuracy of the dynamical METTS algorithm in the limit of low temperatures. Results for spectral functions at different temperatures, including $T=0$ from DMRG, are shown in Fig. 5. We focus on the $16\times 4$ cylinder at $J/J_{D}=0.63$ and two wave vectors $\bm{q}=(\pi,0)$ and $\bm{q}=(\pi/2,\pi/2)$. Our maximal time horizon for the real-time evolution has been chosen as $\Omega/J_{D}=50$. Moreover, we collected $R=100$ independent METTS states for each temperature and observed that this allows sharp error estimates with a large signal-to-noise ratio, as apparent by the small error bars in Fig. 5. The time evolution has been performed using 2TDVP with maximal bond dimensions $\chi_{\max}=1000,1500,2000$. Comparing different bond dimensions for different seeds leads us to conclude that for this particular system, the sample real-time correlation functions $C^{i}(t)$ are already well-converged at $\chi_{\max}=1000$. The spectral function from DMRG has similarly been obtained by a single time-evolution up to time $\Omega/J_{D}=50$. We observe a structured spectrum, where the dominant peak corresponds to single triplon excitations, but also secondary and tertiary peaks corresponding to multi-triplon excitations of the Shastry-Sutherland model. The good convergence towards the DMRG results for $T\rightarrow 0$, where even small features such as secondary and tertiary peaks are well-resolved, demonstrates the accuracy of the approach at low temperatures.
@@ IV.1 Two variants of complex-time correlation functions
Real-time evolution of MPS is computationally expensive due to unbounded entanglement growth for generic wave functions. Imaginary-time correlation functions, on the other hand, can be efficiently computed since for imaginary-time $\tau\rightarrow\infty$ every wavefunction $|\phi\rangle$ will converge to the ground state $|\phi\rangle=|\psi_{0}\rangle$ whenever both wave functions have non-vanishing overlap $|\langle\phi|\psi_{0}\rangle|\neq 0$, i.e.,
@@
There are two natural alternatives when defining complex time correlation functions,
@@
The Hermitian correlator $\mathcal{H}_{AB}(z)$ as in Eq. 44, on the other hand, has only recently been investigated in the context of MPS and impurity models [53] and has several intriguing properties which are discussed Sec. IV.4. We choose the name Hermitian, since whenever $A$ is a Hermitian operator, i.e., $A^{\dagger}=A$, the time-evolved operator $\mathcal{U}(z)^{\dagger}A\mathcal{U}(z)$ is also Hermitian. This property is shared with the real-time correlation function, but not with the analytic correlation function, where in general $\mathcal{A}_{AB}(z)^{\star}\neq\mathcal{A}_{BA}(z)$.
There are computational advantages in evaluating this time-dependent correlation function. For convenience, we again introduce the shorthand notations,
@@
Figure 7: Sketch of the complex contours on which the analytical correlation function $\mathcal{A}(z)$ and the Hermitian correlation functions $\mathcal{H}(z)$ are evaluated. For the analytical correlation function, the maximal imaginary time is chosen to be $\Omega_{\tau}=\beta/2$ in order to allow for an efficient and stable numerical algorithm.
@@ IV.2 Analytic complex-time correlation functions
$\mathcal{A}(z)$ $=\braket{e^{iHz}Ae^{-iHz}B}_{\beta}$
@@
$|v_{i}(z)\rangle=e^{-iHz}B|\psi_{i}\rangle=e^{-iHt}e^{-\tau H}B|\psi_{i}\rangle$ (55)
$|w_{i}(z)\rangle=e^{-iH\overline{z}}|\psi_{i}\rangle=e^{-iHt}e^{+\tau H}|\psi_{i}\rangle,$ (56)
with $z=t-i\tau$ similar to the real-time correlation algorithm Fig. 2. However, this approach is plagued by severe numerical instability when computing the state $|w_{i}(t)\rangle$. Computing a positive exponential $e^{+\tau H}|\psi_{i}\rangle$ is numerically highly unstable and prohibits most practical calculations. This is due to the fact that the initial METTS state is a low-energy state. High-energy contributions are truncated during the time evolution but would need to be resolved accurately when performing positive time evolution causing severe numerical instabilities.
When restricting our time horizon to $\Omega_{\tau}=\beta/2$ we can circumvent ever having to compute positive real exponential in the context of the METTS algorithm. Thus, given we are only interested in simulating complex times with $\Omega_{\tau}<\beta/2$ we propose that the analytic correlator can be evaluated in a numerically stable way. This is done by absorbing part of the positive time evolution into the negative time evolution performed for the METTS states. Notice that a smaller complex angle $\theta$ allows a longer real-time component $\Omega_{t}=\Omega\cos\theta$ while still fulfilling $\Omega_{\tau}=\Omega\sin\theta<\beta/2$.
The key idea is that the positive imaginary time evolution can be absorbed by the negative imaginary time evolution performed when calculating an individual METTS |$\ket{\psi_{i}}$. For this, we rewrite Sec. IV.2 as,
$\mathcal{A}$ $(z)=\braket{e^{izH}Ae^{-izH}B}_{\beta}=$
$=\braket{e^{-i\Omega_{t}H}e^{izH}Ae^{-izH}Be^{i\Omega_{t}H}}_{\beta}$
$=\frac{1}{\mathcal{Z}}\sum_{i}p_{i}\langle\psi_{i}|e^{-i\Omega_{t}H}e^{izH}Ae^{-izH}Be^{i\Omega_{t}H}|\psi_{i}\rangle$
$=\frac{1}{\mathcal{Z}}\sum_{i}p_{i}\;\mathcal{A}^{i}(z),$ (57)
where we defined the sample analytic correlation function as,
$\mathcal{A}^{i}(z)$ $=\langle\psi_{i}|e^{-i\Omega_{t}H}e^{izH}Ae^{-izH}Be^{i\Omega_{t}H}|\psi_{i}\rangle$
Using a discretized $z=k\delta_{z}$ we can rewrite Eq. 59 as,
$\mathcal{A}^{i}(k\delta_{z})=\bra{\phi^{S-k}_{i}}A\ket{\chi_{i}^{k}}.$ (64)
{algorithm}
[t!] Dynamical METTS for analytic correlator $\mathcal{A}(z)$
@@
To evaluate these expectation values for all discretized values of $z=k\delta_{z}$ ($k=0,\ldots,S$), we compute the states |$\ket{\phi_{i}^{k}}$ for $k=0,\ldots,S$ successively by complex time evolution and store all intermediate states to disk for later use. Then we compute the state |$\ket{\chi^{0}_{i}}=B\ket{\psi_{i}^{S}}$ from which we successively compute |$\ket{\chi^{k}_{i}}$ for $k=1,\ldots,S$ and thereby evaluate the desired correlation functions $\mathcal{A}^{i}(z)=\bra{\phi^{S-k}_{i}}A\ket{\chi_{i}^{k}}$. $\mathcal{A}(z)$ is then obtained by averaging over $\mathcal{A}^{i}(z)$. We summarize the described algorithm in Sec. IV.2.
@@
Sec. IV.2 is in several aspects similar to Fig. 2. Both algorithms require computing the METTS |$\ket{\psi_{i}}$ and then time-evolves two further states to compute time-dependent correlators. Moreover, the basic thermal METTS Markov chain is constructed in both cases. A difference in Sec. IV.2 is the necessity of storing the auxiliary states |$\ket{\phi_{i}^{k}}$. On present-day computers, this typically neither imposes severe time nor memory constraints. However, in the limit of small angle $\theta$ the memory requirement can be considerable due to the large number of discrete time steps and requires corresponding storage capacities.